​(Related to Checkpoint​ 5.2)  ​(Compound interest with​ non-annual periods​) You just received a bonus ​$ 5,000.

a.  Calculate the future value of ​$​, given that it will be held in the bank for years and earn an annual interest rate of percent.

b.  Recalculate part ​(a​) using a compounding period that is​ (1) semiannual and​ (2) bimonthly.

c.  Recalculate parts ​(a​) and ​(b​) using an annual interest rate of percent.

d.  Recalculate part ​(a​) using a time horizon of years at an annual interest rate of percent.

e.  What conclusions can you draw when you compare the answers in parts ​(c​) and ​(d​) with the answers in parts ​(a​) and ​(b​)?

.  What is the future value of ​$5,000 in a bank account for 10 years at an annual interest rate of 5 ​percent?

​$8,144.58,144.5  ​(Round to the nearest​ cent.)

b.  What is the future value of ​$5,000 in a bank account for 10 years at 5 percent compounded​ semiannually? ​$8,193.18,193.1  

What is the future value of ​$5,000 in a bank account for 10 years at 5 percent compounded​ bimonthly?
Answer #1

What is the future value of ​$5,000 in a bank account for 10 years at an annual interest rate of 5 ​percent?

​=5000*1.05^10
=8144.47313

b.  What is the future value of ​$5,000 in a bank account for 10 years at 5 percent compounded​ semiannually?

=5000*(1+5%/2)^(2*10)
=8193.08220

What is the future value of ​$5,000 in a bank account for 10 years at 5 percent compounded​ bimonthly?

=5000*(1+5%/6)^(10*6)
=8226.54467